OpenStudy (anonymous):

In the state of Colorado there is a sales tax of 2.9% on all the goods purchased. If Jan buys a television for $599, what will be the final bill, including sales tax?

OpenStudy (campbell_st):

find 2.9% of price and add it to the cost of the TV

OpenStudy (anonymous):

$599*(.029)=14.5 599+14.5=613.5

OpenStudy (anonymous):

oh thanks.

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

uh.. my book said that 616.37 is the answer.

OpenStudy (anonymous):

599 * 0.029 = 17.371 not 14.5

OpenStudy (anonymous):

yea i got the same thing as your book I just accidentally typed it in my calculator wrong

OpenStudy (anonymous):

the work above is correct!

OpenStudy (anonymous):

lol thx.

OpenStudy (anonymous):

np! :)